OpenStudy (anonymous):

pre-cal f(g(-2)) f(x)=3x^2+4 g(x)=5x I did this: g(-2)= 5(-2) g(-2)= -10 f(g(-2)): 3(-10)^2+4 f(g(-2))=-296 But the correct answer is 304. Where did I make my mistake? Thank you for any help!!

OpenStudy (anonymous):

3(-10)^2+4=3*100+4

OpenStudy (anonymous):

This was what I got. But my calculator got -296. Does the (-10)^2 cancel out the negative?

OpenStudy (anonymous):

(-10)*(-10)=100

OpenStudy (anonymous):

n calculator U need put -10 in ()

OpenStudy (anonymous):

Ohh! Ok, tysm!!

OpenStudy (anonymous):

:-))) thanks for medal